OpenStudy (anonymous):

Simplify the following. (2x/(x^2-6x+9)) - (1/(x+1)) - (8/(x^2-2x-3))

OpenStudy (anonymous):

(2 x)/(x^2-6 x+9)-8/(x^2-2 x-3)-1/(x+1) Maybe that will help =)

OpenStudy (anonymous):

I'm so lost... Could you explain a little more? I knew this last year but I've forgotten over the summer. -_-

OpenStudy (anonymous):

It looks like this if this helps: \[\frac{ 2x }{ x^2-6x+9} - \frac{ 1 }{ x+1 } - \frac{ 8 }{ x^2-2x-3 }\]

OpenStudy (anonymous):

ok so wouldn't you cross divide any similar factors like

OpenStudy (anonymous):

..... at least that's what I think

OpenStudy (anonymous):

you can't do that... the crossing out thing only works if you have factors on the top and the bottom. It still has substraction symbols outside of actual factors.

OpenStudy (anonymous):

huh.... wait i don't get it 2x goes into 6x 3 times so why wouldn't it be 3x? sorry i feel like i'm confusing you not helping you.....

OpenStudy (anonymous):

No, it's okay. The reason you can't do that is because you can't pick and choose from the polynomial what you want to simplify. You have to factor the top and the bottom first to do that. For instance, \[\frac{ x^2+4x+3 }{ x^2+2x+1 }\]\[\frac{ (x+3)(x+1) }{ (x+2)(x+1) }\] I can't suddenly cancel the 4 and make it a two and make the 2 on the bottom a 1. It doesn't work. I have to factor both things first because in order to divide I need numbers that multiply, like this: \[\frac{ (x+3)(x+1) }{ (x+2)(x+1) }\] Now you can cancel the two (x+1) 's because they are multiplying the stuff around them, not subtracting or adding.

OpenStudy (anonymous):

But seriously, I really need help on this one. It's driving me crazy.

OpenStudy (anonymous):

I get it!.... but didn't you kinda solve the question you have? :)

OpenStudy (anonymous):

I wish.... :/

OpenStudy (anonymous):

lol give me a minute... let's try this again shall we? :)

OpenStudy (anonymous):

All right... it's just that I know there was a technique I learned but I don't remember anything!

OpenStudy (anonymous):

Plus, I bumped the question and it's not going to the top. -_-

OpenStudy (anonymous):

it's fine i hate when that happens though it's like i know it i know it i know it.... I DON'T KNOW IT... and i know the first time i had a question and tried to bump it up top it wouldn't work and i was like greattttttt..... LOL :)

OpenStudy (anonymous):

ok so i'm not sure if i did this right but here goes so the end result would be x-8/-8+6........ what do you think?